OpenStudy (anonymous):

Given the following unbalanced chemical equation, how many moles of nitrogen monoxide are needed to react with excess ammonia to produce 32.6 grams of water? ___NH3 (g) + ___ NO (g) arrow ___N2 (g) + ___H2O (l)

OpenStudy (abhisar):

@rosieclow \(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (abhisar):

Can u balance the given equation ?

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

pretty sure when its balanced it should be 4 NH3 (g) + 6 NO (g) = 5 N2 (g) + 6 H2O (l)

OpenStudy (abhisar):

CORRECT ! \(\large\sf 4 NH_3 (g) + 6 NO (g) = 5 N_2 (g) + 6 H_2O (l)\)

OpenStudy (anonymous):

so how many moles of nitrogen monoxide are needed to react with excess ammonia to produce 32.6 grams of water?

OpenStudy (anonymous):

yes that sounds right!

OpenStudy (anonymous):

A. 5.53 x 10-1 moles NO B. 10.9 moles NO C. 32.6 moles NO D. 1.81 moles NO these are my options by the way

OpenStudy (abhisar):

6 moles of NO is required for creating 6 moles of water i.e 180 of NO is required for creating 108 grams of water.

OpenStudy (abhisar):

So, \(\huge\sf \frac{180}{108}\times32.6\) grams of NO will be required for creating 32.6 grams of water. This will be equal to 54.33 grams of NO

OpenStudy (abhisar):

Now can u temme the molar mass of NO ?

OpenStudy (anonymous):

I can't :(

OpenStudy (abhisar):

Temme the atomic mass of N and O

OpenStudy (anonymous):

29.9

OpenStudy (abhisar):

yes that should be the molecular mass of NO. Now to get the moles of NO simply divide 54.33 grams with 30. This will give you the answer !

OpenStudy (anonymous):

so its 1.8 or D! thank you Abhisar you're the best

OpenStudy (abhisar):

yes it's D

OpenStudy (abhisar):

Did you know you can give medals to people who help you with questions? Just click the \(\color{blue}{\text{blue "best response" button}}\) to show appreciation for the people who help you.

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)